# Physics of droplet regulation in biological cells

### v2

@@ 3.3 Complex interfaces
Many proteins molecules are amphiphilic, and can thus act as surfactants, preferentially associating to the interface [263, 264, 265] to form Pickering emulsions. In general, surfactants stabilise droplet interfaces, resisting both dissolution and coalescence. The formation of complex interfaces via preferential protein association can therefore improve structural integrity of condensates and assist in size regulation [266], as well as leading to complex droplet morphologies [267]. Note that although surfactants usually result in kinetic droplet stabilisation, some complex particles may also result in thermodynamic stabilisation [268, 269].
The amphiphilic properties of protein molecules depend on their particular conformational state. Correspondingly, the chemical switching of a particular protein from a surfactant to liquid-like form has been implicated in chromosome clustering during mitotic exit [270]. Further, interfaces themselves can alter protein conformations, typically leading to extended molecular configurations [271]. Variations in either surfactant configuration or compositional state along an interface can give rise to surface tension gradients. These gradients drive Marangoni flows at the surface of the droplet, which may then lead to droplet swimming [272, 64, 273], thus highlighting how complex interfacial properties are able to affect the behaviour of the whole droplet.
@@ 3.4 Electrostatic effects
Charge is another important biomolecular property that is involved in the regulation of condensates [274]. On the one hand, charge patterns on protein surfaces can mediate specific, short-ranged interactions [275]. On the other hand, a net charge induces electrostatic interactions that can potentially range over long distances. Indeed, many condensates contain charged components [276, 277], which affects interfacial properties [278, 279, 262, 280], slows down coarsening [281, 282], limits droplet size [155, 283, 284], influences droplet–environment interactions [285, 286], and can potentially lead to droplet division [287, 288]. The dynamics of charged species can be described by the classical electrostatic framework. The central quantity is the charge density, $\varrho(\boldsymbol{r})=\sum_{i}q_{i}\phi_{i}$, which depends on the volume fractions $\Phi=\bigl{(}\phi_{1},\ldots,\phi_{{N_{\mathrm{c}}}}\bigr{)}$ of all species $i=1,\ldots,{N_{\mathrm{c}}}$, each with charge $q_{i}$ per unit volume. The charge density $\varrho$ induces an electrostatic potential $\varphi(\boldsymbol{r})$, which satisfies the Poisson equation
@@
Over short distances, electrostatic effects mediate attraction between oppositely charged molecules, which can lead to phase separation from a neutral solvent via complex coacervation [291, 25, 292, 293]. Coacervates have been studied for about a century [294] and have been implicated in the origin of life [295, 296]. In the simplest case, the attraction of oppositely charged molecules can be captured by an additional term in the local free energy $f$ given by equation (2.1). For instance, in $d=3$ dimensions, Voorn–Overbeek theory leads to the free energy density [297, 298]
@@
where the non-dimensional factor $\alpha=\frac{1}{3}(4\pi)^{2}\ell_{\mathrm{B}}^{3/2}\nu_{0}^{-1/2}$ governs the strength of the electrostatic interaction. Here, $\ell_{\mathrm{B}}=e^{2}/(4\pi\varepsilon k_{\mathrm{B}}T)$ is the Bjerrum length, which describes the distance beyond which thermal energy dominates electrostatic repulsion of two elementary charges $e$ [299]. Since the permittivity $\varepsilon$ of cytosol is comparable to water [300], a typical Bjerrum length is $\ell_{\mathrm{B}}\approx 0.7\,\mathrm{nm}$. We thus have $\alpha\sim 0.1\ldots 1$ for typical molecular volumes of $\nu_{0}\sim 10^{3}\ldots 10^{5}\,\mathrm{nm^{3}}$, indicating that electrostatic effects can lower the free energy significantly.
@@ 3.5 Multiphase coexistence
Cells often comprise many different types of condensates, consistent with Gibbs phase rule, which limits the number of different phases to be smaller than or equal to the number of different components; see section 2.1.1. The number of coexisting phases and their compositions can generally be determined by minimising the free energy of the entire system, which leads to the coexistence conditions given by equations (2.5). However, even for the simple Flory–Huggins theory given by equation (2.1), these relations are difficult to solve since they are non-linear and depend on all interaction parameters $\chi_{ij}$ and compositions $\bar{\phi}_{i}$. It is thus not surprising that the resulting phase diagrams, showing the equilibrium phases as a function of the system composition, can be complex [103]. Even in the simplest case of two interacting solutes, the topology of the phase diagram depends on the interaction strength; see Figure 3.1. This multi-component, multi-phase problem has been tackled by linear stability analysis of the homogeneous state [301, 302, 303, 118, 304], grand-canonical simulations [305, 306], continuation methods [307], convexification of the free energy density [103], and machine learning methods [308].
3.5.1  Random mixtures
@@
The behaviour of a mixture with random interactions also depends on its overall composition. While it is convenient to analyse equimolar mixtures, where $\phi_{i}=({N_{\mathrm{c}}}+1)^{-1}$, variations in composition strongly affect the instability of the homogeneous state [304]. Such variations can be captured by sampling a symmetric Dirichlet distribution with the density function
$P(\{\phi_{0},\ldots,\phi_{N_{\mathrm{c}}}\};\alpha)=\frac{\Gamma(\alpha{N_{\mathrm{c}}}+\alpha)}{\Gamma(\alpha)^{{N_{\mathrm{c}}}+1}}\prod_{i=0}^{N_{\mathrm{c}}}\phi_{i}^{\alpha-1}\;,$ (3.7)
where $\alpha>0$ is a shape parameter. The equimolar case is recovered in the limit $\alpha\rightarrow\infty$; $\alpha=1$ samples all compositions uniformly; and $\alpha<1$ result in mixtures with a few dominating species. Taken together, this protocol enables study of the average behaviour of a multi-component liquid with random interactions.
3.5.2  Instability of homogeneous states
Phase separation of a multi-component mixture must occur when the homogeneous state is unstable, which is the case when the Hessian $H_{ij}=\partial^{2}f/\partial\phi_{i}\partial\phi_{k}$ exhibits at least one negative eigenvalue. If there are multiple negative eigenvalues, the eigenvectors generally point in different directions, suggesting that the system segregates into various phases. The number of negative eigenvalues, $N_{\mathrm{u}}$, can be predicted for the Flory–Huggins model, for which
$H_{ij}$ $=\frac{k_{\mathrm{B}}T}{\nu_{0}}\left(\frac{\delta_{ij}}{\Omega_{i}\phi_{i}}+\frac{1}{\Omega_{0}\phi_{0}}+\chi_{ij}\right)\;,$ (3.8)
following from equation (2.1). Using the random interactions and compositions introduced in section 3.5.1, random matrix theory can be used to predict $N_{\mathrm{u}}$ [309, 301, 118]. For a large number of components and uniformly distributed compositions ($\alpha=1$), the average value of $N_{\mathrm{u}}$ is approximately [310]
@@ 3.5.4  Towards realistic systems
The scaling result given in equation (3.10) predicts some features of condensate coexistence observed in cells, but is likely not quantitatively accurate. This discrepancy is because the cell’s composition is clearly not random, but instead homeostatically controlled, and interactions between biomolecules are also not random, but structured [312, 206, 313, 314, 315] and also under cellular control [316, 97, 317, 318, 319]. Moreover, interactions are typically complex [320], likely involving higher-order interactions [321] and the effect of polymer dynamics [322, 212, 210]. Note also that the number ${N_{\mathrm{p}}}$ of coexisting phases is only the most basic descripter of multi-component phase separation. The compositions of different phases are also crucial for cellular function since they give rise to the droplets’ identities. Along these lines, it will also be interesting to understand what determines the interfacial parameters $\kappa_{ij}$ in equation (2.6), since these are crucial for predicting surface tensions and thus multiphase morphologies [323, 103, 324]. In the simplest case, surface tension and other related quantities might simply scale as a power law of the distance to a critical point [325], but whether this situation holds generally is unclear. Besides surface tension, droplet morphology is also shaped by other objects in the surrounding environment, as discussed in section 4 below.
@@ 4.1 Theory of wetting
The interaction between a multi-component mixture and external structures can be captured by including a surface energy term in the free energy, which acts at the boundary of the mixture. In the continuous description presented in section 2.1, we thus have a total energy given by
@@ 4.1.1  Partial, full and pre- wetting
To simplify the discussion, we focus on a binary fluid mixture in contact with a single external boundary (be this a rigid solid, soft solid, or liquid). We then extend the effective droplet model introduced in section 2.2, in which a dense (droplet) phase of volume $V$ and composition $\phi^{\mathrm{eq}}_{\mathrm{in}}$, coexists with a dilute phase of composition $\phi^{\mathrm{eq}}_{\mathrm{out}}$. The energetic interactions of the binary mixture with the boundary are described by surface energies $\gamma_{\mathrm{in}}=f_{\mathrm{surf}}(\phi^{\mathrm{eq}}_{\mathrm{in}})$ and $\gamma_{\mathrm{out}}=f_{\mathrm{surf}}(\phi^{\mathrm{eq}}_{\mathrm{out}})$, for the dense and dilute phases, respectively. Assuming that the boundary has a fixed surface area $S_{0}$, the total free energy of the mixture is then
@@
In undersaturated systems ($\bar{\phi}<\phi^{(0)}_{\mathrm{out}}$), ${\Delta\!f<0}$, and so isolated droplets cannot form. However, if $\gamma_{\mathrm{in}}<\gamma_{\mathrm{out}}$, and $\gamma_{\mathrm{out}}-\gamma_{\mathrm{in}}$ is sufficiently large, a layer of droplet material can be stabilised along the external surface (figure 4.1C). This phenomenon is known as pre-wetting. As the concentration of the homogenous phase approaches the saturation concentration, the thickness of the pre-wetted layer diverges as ${h\sim-\log(\phi^{(0)}_{\mathrm{out}}-\bar{\phi})}$  [328]. Typically, pre-wetting is only possible in a very narrow portion of the parameter space, but recent work has shown that binding interactions between droplet molecules and the surface can significantly enhance the size of the pre-wetted regime [327, 329]
In a supersaturated system ($\bar{\phi}>\phi^{(0)}_{\mathrm{out}}$), $\Delta\!f>0$, and so macroscopic droplets are able to form regardless of the presence of external boundaries. Nevertheless, surface interactions still have a significant impact on droplet regulation, both in terms of droplet nucleation (see section 4.1.4 below) and droplet morphology. In particular, if $\gamma_{\mathrm{in}}<\gamma_{\mathrm{out}}$ and $\gamma<\gamma_{\mathrm{out}}-\gamma_{\mathrm{in}}$, the dense phase will completely ‘wet’ the surface, forming a macroscopic layer between the surface and the dilute phase. This configuration is known as perfect, or full, wetting. For systems in which $\gamma>|\gamma_{\mathrm{out}}-\gamma_{\mathrm{in}}|$, the dense phase will instead partially wet the surface to form spatially compact droplets. The crossover between these two cases is known as the wetting transition [328]. The different phenomenologies of pre-, partial, and full wetting are highlighted in figures 4.1A-D. Figure 4.1E shows a phase diagram demarcating the transitions between the different wetting regimes.
@@ 4.1.2  Solid substrates
Note that when $\gamma_{\mathrm{out}}>\gamma_{\mathrm{in}}$, $20<\vartheta<\pi/$ 2 and the surface is preferentially wetted by the droplet phase. When $\gamma_{\mathrm{out}}<\gamma_{\mathrm{in}}$, $\pi/2<\vartheta<\pi$ and the surface is preferentially wetted by the dilute phase.
Capillary forces resulting from differential wetting may deform sufficiently soft solid substrates. The elasto-capillary length $\ell_{\textrm{cap}}=\gamma/E$ characterises the relative strength of capillary interactions and elastic deformation, for a solid with stretching modulus $E$. For very small $\ell_{\textrm{cap}}$, resistance to stretching dominates capillary forces, and the substrate is effectively rigid. If $\ell_{\textrm{cap}}$ is comparable to the interfacial width $0pt$ of the droplet, the surface will deform at the contact line, forming small ridges which point towards the droplet [330]. Such deformations result in a modified contact angle described by a generalised Neumann’s law [71],
$\boldsymbol{T}\cdot\boldsymbol{n}-\boldsymbol{\sigma}\cdot\boldsymbol{n}=\frac{\text{d}\Gamma_{\textrm{s}}}{\text{d}s}\boldsymbol{t}\;,$ (4.7)
where $\boldsymbol{T}$ and $\boldsymbol{\sigma}$ are the respective stresses in the solid substrate and droplet phase, $\boldsymbol{n}$ and $\boldsymbol{t}$ are vectors perpendicular and parallel to the surface, and $s$ is a curvilinear coordinate which follows the surface of the substrate. $\Gamma_{\mathrm{s}}$ is the solid surface stress, which may in general depend on the surface strain (Shuttleworth effect [331]). For very soft materials, $\ell_{\textrm{cap}}$ is comparable to droplet radius $R$, and hence the rigidity of the solid is negligible compared to surface tension. In this case, the substrate deforms on the scale of the entire droplet, which thus acquires a lens-like shape [71].
@@ 4.1.3  Interactions with other droplets
If multiple immiscible droplets are present in a given system, they may come into contact, and thus interact. The morphology of the droplet–droplet interaction is governed by the relative strengths of surface tension between the different phases (figure 4.1F). To describe the interaction between two immiscible droplets, we denote the surface tension between the droplets by $\gamma_{12}$, and the respective surface tensions between each droplet and the dilute phase as $\gamma_{1}$ and $\gamma_{2}$. If $\gamma_{12}\gg\gamma_{1}\approx\gamma_{2}$, then droplet–droplet interfaces are relatively energetically costly, and so the droplets will remain separated. If $\gamma_{1}\approx\gamma_{2}\approx\gamma_{12}$, the droplets will partially wet each other. The contact angle of the droplet–droplet interface is then determined by Neumann’s law, which can be obtained by setting $\textrm{d}\Gamma_{\textrm{s}}/\textrm{d}s=0$ in equation (4.7). Finally, if either $\gamma_{1}$ or $\gamma_{2}$ is sufficiently large compared to the other two surface tensions, perfect wetting conditions may occur, resulting in the engulfment of the droplet with large surface tension by the other. Such ‘double emulsion’ structures have been observed in several sub-cellular structures, including nucleoli and stress granules [332].
@@ 4.1.4  Heterogeneous nucleation
$\Delta F=\dfrac{4\pi\gamma^{3}}{3\Delta\!f^{2}}(2+\cos\vartheta)(1-\cos\vartheta)^{2}\;.$ (4.8)
For perfectly non-wetting conditions ($\vartheta=\pi$) the classic homogenous nucleation result (equation 2.34) is recovered. Any other contact angle results in preferential nucleation at the surface. Analogous expressions can also obtained for more complex surface shapes [335]. By tuning the strength of interactions between droplet material and other cellular structures, cells are thus able to regulate the location and rate at which condensates are nucleated [336, 337, 338, 339].

### v3

@@ 3.3 Complex interfaces
Many proteins molecules are amphiphilic, and can thus act as surfactants, preferentially associating to the interface [283, 284, 285, 286]. In general, surfactants stabilise droplet interfaces, resisting both dissolution and coalescence. The formation of complex interfaces via preferential protein association can therefore improve structural integrity of condensates and assist in size regulation [287], as well as leading to dynamical anisotropies at the interface [288]. Note that although surfactants usually result in kinetic droplet stabilisation, some complex particles may also result in thermodynamic stabilisation to form Pickering emulsions [161, 289, 290].
The amphiphilic properties of protein molecules depend on their particular conformational state [291]. Correspondingly, the chemical switching of a particular protein from a surfactant to liquid-like form has been implicated in chromosome clustering during mitotic exit [292]. Further, interfaces themselves can alter protein conformations, typically leading to extended molecular configurations [293]. Variations in either surfactant configuration or compositional state along an interface can give rise to surface tension gradients. These gradients drive Marangoni flows at the surface of the droplet, which may then lead to droplet swimming [294, 68, 295], thus highlighting how complex interfacial properties are able to affect the behaviour of the whole droplet.
@@ 3.4 Electrostatic effects
Charge is another important biomolecular property that is involved in the regulation of condensates [296]. On the one hand, charge patterns on protein surfaces can mediate specific, short-ranged interactions [297]. On the other hand, a net charge induces electrostatic interactions that can potentially range over long distances. Indeed, many condensates contain charged components [298, 299], which affects interfacial properties [300, 301, 302, 282, 303], slows down coarsening [181, 304], limits droplet size [162, 305, 306], influences droplet–environment interactions [307, 308], and can potentially lead to droplet division [309, 310]. The dynamics of charged species can be described by the classical electrostatic framework [311, 312, 313]. The central quantity is the charge density, $\varrho(\boldsymbol{r})=\sum_{i}q_{i}\phi_{i}$, which depends on the volume fractions $\Phi=\bigl{(}\phi_{1},\ldots,\phi_{{N_{\mathrm{c}}}}\bigr{)}$ of all species $i=1,\ldots,{N_{\mathrm{c}}}$, each with charge $q_{i}$ per unit volume. The charge density $\varrho$ induces an electrostatic potential $\varphi(\boldsymbol{r})$, which satisfies the Poisson equation
@@
Over short distances, electrostatic effects mediate attraction between oppositely charged molecules, which can lead to their phase separation from a neutral solvent [316, 26, 317, 318]. This complex coacervation has been studied for about a century [319] and has been implicated in the origin of life [320, 321]. In the simplest case, the attraction of oppositely charged molecules can be captured by an additional term in the local free energy $f$ given by equation (2.1). For instance, in $d=3$ dimensions, Voorn–Overbeek theory leads to the free energy density [322, 323]
@@
where the non-dimensional factor $\alpha=\frac{1}{3}(4\pi)^{2}\ell_{\mathrm{B}}^{3/2}\nu_{0}^{-1/2}$ governs the strength of electrostatic interactions. Here, the Bjerrum length $\ell_{\mathrm{B}}=e^{2}/(4\pi\varepsilon k_{\mathrm{B}}T)$ describes the distance beyond which thermal energy dominates electrostatic repulsion of two elementary charges $e$ [324]. A typical value is $\ell_{\mathrm{B}}\approx 0.7\,\mathrm{nm}$ since the permittivity $\varepsilon$ of cytosol is comparable to water [325]. We thus have $\alpha\sim 0.1\ldots 1$ for typical molecular volumes of $\nu_{0}\sim 10^{3}\ldots 10^{5}\,\mathrm{nm^{3}}$, and the electrostatic contribution becomes comparable to the enthalpic contributions in $f$ if $|q_{i}|\sim e$, indicating that electrostatic effects can lower the free energy significantly. The theory given by equation (3.6) qualitatively explains complex coacervation, but it cannot predict realistic coacervates quantitatively [326], particularly when the oppositely charged molecules are asymmetric [327, 304, 162, 305, 306, 328].
@@ 3.5 Multiphase coexistence
Cells often comprise many different types of condensates, consistent with Gibbs phase rule, which limits the number of different phases to be smaller than or equal to the number of different components; see section 2.1.1. The number of coexisting phases and their compositions can generally be determined by minimising the free energy of the entire system, which leads to the coexistence conditions given by equations (2.5). The resulting phase diagrams, showing equilibrium phases as a function of the system’s composition, can be complex [108]. Even in the simplest case of two interacting solutes, the topology of the phase diagram depends on the interaction strength; see Figure 3.1. This multi-component, multi-phase problem has been tackled by linear stability analysis of the homogeneous state [329, 330, 331, 124, 218], grand-canonical simulations [332, 333], continuation methods [334], convexification of the free energy density [335, 336, 108], and machine learning methods [337].
3.5.1 Instability of homogeneous states
Phase separation of a multi-component mixture must occur when the homogeneous state is unstable, which is the case when the Hessian $H_{ij}=\partial^{2}f/\partial\phi_{i}\partial\phi_{k}$ exhibits at least one negative eigenvalue. If there are multiple negative eigenvalues, the eigenvectors generally point in different directions, suggesting that the system segregates into various phases. The number of negative eigenvalues, $N_{\mathrm{u}}$, can be predicted for the Flory–Huggins model given by equation (2.1), for which
$H_{ij}$ $=\frac{k_{\mathrm{B}}T}{\nu_{0}}\left(\frac{\delta_{ij}}{\Omega_{i}\phi_{i}}+\frac{1}{\Omega_{0}\phi_{0}}+\chi_{ij}\right)\;.$ (3.7)
@@ 3.5.3 Towards realistic systems
The scaling result given in equation (3.9) predicts some features of condensate coexistence observed in cells, but is likely not quantitatively accurate. This discrepancy is because the cell’s composition is clearly not random, but instead homeostatically controlled, and interactions between biomolecules are also not random, but structured [341, 342, 221, 343, 344, 345] and also under cellular control [346, 102, 347, 348, 349]. Moreover, interactions are typically complex [350], likely involving higher-order interactions [351] and the effect of polymer dynamics [352, 227, 225]. Note also that the number $N_{\mathrm{p}}$} of coexisting phases is only the most basic descripter of multi-component phase separation. The compositions of different phases are also crucial for cellular function since they give rise to the droplets’ identities. Along these lines, it will also be interesting to understand what determines the interfacial parameters $\kappa_{ij}$ in equation (2.6), since these are crucial for predicting surface tensions and thus multiphase morphologies [353, 354, 355, 356, 108, 357]. In the simplest case, surface tension and other related quantities might simply scale as a power law of the distance to a critical point [358], but whether this situation holds generally is unclear. Besides surface tension, droplet morphology is also shaped by other objects in the surrounding environment, as discussed in section 4 below.
@@ 4.1 Theory of wetting
The interaction between a multi-component mixture and external structures can be captured by including a surface energy term in the free energy, which acts at the boundary of the mixture. In the continuous description presented in section 2.1, we thus have a total energy given by [360, 361]
@@ 4.1.1 Partial, full and pre- wetting
To simplify the discussion, we focus on a binary fluid mixture in contact with a single external boundary (be this a rigid solid, soft solid, or liquid). We then extend the effective droplet model introduced in section 2.2, in which a dense (droplet) phase of volume $V$ and composition $\phi^{\mathrm{eq}}_{\mathrm{in}}$, coexists with a dilute phase of composition $\phi^{\mathrm{eq}}_{\mathrm{out}}$. The energetic interactions of the binary mixture with the boundary are described by surface energies $\gamma_{\mathrm{in}}=f_{\mathrm{surf}}(\phi^{\mathrm{eq}}_{\mathrm{in}})$ and $\gamma_{\mathrm{out}}=f_{\mathrm{surf}}(\phi^{\mathrm{eq}}_{\mathrm{out}})$, for the dense and dilute phases, respectively. Assuming that the boundary has a fixed surface area $S_{0}$, the total free energy of the mixture is then [362, 363]
@@
[…]
@@ 4.1.2 Solid substrates
Note that when $\gamma_{\mathrm{out}}>\gamma_{\mathrm{in}}$, $0<\vartheta<\pi/2$ and the surface is preferentially wetted by the droplet phase. When $\gamma_{\mathrm{out}}<\gamma_{\mathrm{in}}$, $\pi/2<\vartheta<\pi$ and the surface is preferentially wetted by the dilute phase.
Capillary forces resulting from differential wetting may deform sufficiently soft solid substrates. The elasto-capillary length $\ell_{\textrm{cap}}=\gamma/E$ characterises the relative strength of capillary interactions and elastic deformation, for a solid with stretching modulus $E$. For very small $\ell_{\textrm{cap}}$, resistance to stretching dominates capillary forces, and the substrate is effectively rigid. If $\ell_{\textrm{cap}}$ is comparable to the interfacial width $0pt$ of the droplet, the surface will deform at the contact line, forming small ridges which point towards the droplet [367, 75]. For very soft materials, $\ell_{\textrm{cap}}$ is comparable to droplet radius $R$, and hence the rigidity of the solid is negligible compared to surface tension. In this case, the substrate deforms on the scale of the entire droplet, which thus acquires a lens-like shape [75], with angles between each interface determined by Neumann’s law $\gamma\boldsymbol{t}+\gamma_{\mathrm{in}}\boldsymbol{t}_{\textrm{in}}+\gamma_{\mathrm{out}}\boldsymbol{t}_{\textrm{in}}=0$}, where $\boldsymbol{t}$, $\boldsymbol{t}_{\textrm{in}}$, and $\boldsymbol{t}_{\textrm{out}}$ are the tangential unit vectors along the droplet–dilute, droplet–substrate, and dilute–substrate interfaces, respectively.
@@ 4.1.3 Interactions with other droplets
If multiple immiscible droplets are present in a given system, they may come into contact, and thus interact. The morphology of the droplet–droplet interaction is governed by the relative strengths of surface tension between the different phases (figure 4.1F), and may be regulated by surfactant proteins that associate to interfaces and modify these surface tensions [368]. To describe the interaction between two immiscible droplets, we denote the surface tension between the droplets by $\gamma_{12}$, and the respective surface tensions between each droplet and the dilute phase as $\gamma_{1}$ and $\gamma_{2}$. If $\gamma_{12}\gg\gamma_{1}\approx\gamma_{2}$, then droplet–droplet interfaces are relatively energetically costly, and so the droplets will remain separated. If $\gamma_{1}\approx\gamma_{2}\approx\gamma_{12}$, the droplets will partially wet each other. The contact angle of the droplet–droplet interface is then determined by Neumann’s law, analogously to droplets wetting a very soft surface as described above. Finally, if either $\gamma_{1}$ or $\gamma_{2}$ is sufficiently large compared to the other two surface tensions, perfect wetting conditions may occur, resulting in the engulfment of the droplet with large surface tension by the other. Such ‘double emulsion’ structures have been observed in several sub-cellular structures, including nucleoli and stress granules [369].
@@ 4.1.4 Heterogeneous nucleation
$\Delta\!F_{\mathrm{nucl}}=\dfrac{4\pi\gamma^{3}}{3\Delta\!f^{2}}(2+\cos\vartheta)(1-\cos\vartheta)^{2}\;.$ (4.7)
This expression is obtained analogously to the energy barrier for homogeneous nucleation (equation 2.34) by finding the maximum value of $\Delta F$ prescribed by equation (4.4). For perfectly non-wetting conditions ($\vartheta=\pi$) the classic homogeneous nucleation result is recovered. Any other contact angle results in preferential nucleation at the surface. Analogous expressions can also obtained for more complex surface shapes [372]. By tuning the strength of interactions between droplet material and other cellular structures, cells are thus able to regulate the location and rate at which condensates are nucleated [373, 374, 375, 376].